Economic Life

The Scampini Supplies Company recently purchased a new delivery truck. The new truck cost $22,500, and it is expected to generate net after-tax operating cash flows, including depreciation, of $6,250 per year. The truck has a 5-year expected life. The expected salvage values after tax adjustments for the truck are given below. The company's cost of capital is 11 percent.

Year Annual Operating Cash Flow Salvage Value
0 -$22,500 $22,500
1 6,250 17,500
2 6,250 14,000
3 6,250 11,000
4 6,250 5,000
5 6,250 0
  1. What is the optimal number of years to operate the truck?
Answer #1
Present Value (PV) of Cash Flow:
(Cash Flow)/((1+i)^N)
i=Discount Rate=Cost of Capital=11%
N=Year of Cash Flow
N A B=A/(1.11^N) C=-(22500-B) D E=D/(1.1^N) F G=C+F H(Using PMT function with Rate=11%,Nper=N,PV=-G)
End of year Salvage Value PV of Salvage Value PV of net capital cost Annual Operating Cash Flow PV of annual Cash Flow Cumulative PV of annual Cashflow PV of total Cash Flow Equivalent annualcost
0
1 $17,500 $15,766 ($6,734) $6,250 $         5,682 $        5,682 ($1,052) ($1,168) (Using PMT function with Rate=11%,Nper=1,PV=-1052)
2 $14,000 $11,363 ($11,137) $6,250 $         5,165 $     10,847 ($290) ($169) (Using PMT function with Rate=11%,Nper=2,PV=290)
3 $11,000 $8,043 ($14,457) $6,250 $         4,696 $     15,543 $1,086 $444 (Using PMT function with Rate=11%,Nper=3,PV=-G1086)
4 $5,000 $3,294 ($19,206) $6,250 $         4,269 $     19,812 $605 $195 (Using PMT function with Rate=11%,Nper=4,PV=-605)
5 $0 $0 ($22,500) $6,250 $         3,881 $     23,692 $1,192 $323 (Using PMT function with Rate=11%,Nper=4,PV=-1192)
Annual EquivalentCash Flow is HIGHEST for 3 years $444
The Truck should be operated for OPTIMUM 3 years
Editing Number Clipboard Font Alignment Styles Cells =PMT(11 %, G12,-N12) fix O12 G H K L M N Present Value (PV) of Cash Flow